How do I put x^2+20x+103 in vertex form?
Answer by richwmiller(17219)

parabola
focus | (-10, 13/4)~~(-10, 3.25)
vertex | (-10, 3)
semi-axis length | 1/4
focal parameter | 1/2
eccentricity | 1
directrix | y = 11/4
y=x^2+20x+103
y= a(x-h)2+k
(h,k) is the vertex
y=(x+10)^2+3